OpenStudy (anonymous):

What is the simplified form of the expression? 1/3(21m + 27) A. 63m + 9 B. 7m + 9 C. 7m + 81 D. 7m + 27 im thinking D

OpenStudy (anonymous):

Alright so we have to find the steps. 1/3(21m+27) =(1/3)(21m+27) =(1/3)(21m)+(1/3)(27) =7m+9 7m+9 is your final answer, B.

OpenStudy (anonymous):

how??

OpenStudy (damonte212):

(1/3)(21m+27) = 21m/3 + 27/3 = 7m + 9 <-------- answer